### v1

@@ 4.1 Method: Description and evaluation of the experiment
This experiment considers the primary use case of the dataset: to extract a set of relevant opinions from individuals based on the transcript of a public hearing.
As shown in section 3.3, the transcripts of PublicHearingBR are extensive texts. Depending on the model, the text can exceed its context window limit. Due to this limitation, we opted to test ChatGPT because it is more capable of handling larger inputs. This also simulates the perspective of an end user of the platform and it allows us to estabilish a baseline for the dataset.
We created a custom GPT with a specialized instruction prompt (Figure 4) that provides detailed instructions on how to read and interpret the transcripts. It shows the format of the input data and instructs the language model to extract, from the transcript, a structured summary containing relevant opinions of the individuals.
Once the structured summaries for all samples have been extracted by ChatGPT, it is necessary to compare them with the ground truth available in the PublicHearingBR dataset to assess their quality.
Since this form of summarization is hybrid, using the ROUGE metric [7] to evaluate the system is inadequate, as it is based on n-grams rather than the text’s semantics. We propose that the evaluation covers three aspects: 1. recall: to determine the percentage of relevant opinions that the experiment was able to return; 2. precision: to check the percentage of returned opinions that are relevant; 3. hallucination: to ensure that the extracted opinions are present in the transcript, even if they are not in the dataset.
@@ 4.1.1 Recall and precision
We use the following three steps to calculate recall and precision:
@@
    The first step involves mapping the individuals mentioned in the transcript to those in the metadata available in PublicHearingBR. Since names might differ (due to abbreviations or titles, for example), a language model is initially asked to perform this mapping. This was done using GPT-4o with the prompt shown in Figure 5.
@@
    The second step involves mapping, for each individual identified in the previous step, the opinions extracted from the transcript that are in the PublicHearingBR dataset. This is necessary because, as these are natural language texts, it is important to analyze whether the opinions are semantically the same. This mapping was also performed using GPT-4o with the prompt shown in Figure 6.
@@
    Once the mapping is complete, it is possible to verify the number of opinions from the dataset that were returned and, thereby, calculate recall and precision.
In this case, the recall is a relevant metric, as it informs if the opinions returned by human evaluators are also returned by the summarization system. It is worth noting that the recall calculated through the mapping done by the LLM is the essence of the Check-Eval method. The main difference is that, in Check-Eval, a reference checklist is automatically generated by an LLM from a text [14], while in this method, the checklist, which in this case consists of the opinions of the individuals, the reference checklist is obtained directly from the dataset.
Low precision is not necessarily a bad outcome. When writing an article, a journalist may choose to omit other relevant opinions due to space constraints, for example. However, it is important to evaluate whether the summarization system is hallucinating. This is critical because the goal is to write a news article (or provide inputs for that), where false information is unacceptable. Besides, as the public audience is a debate, with participants sharing their ideas and opinions, it is essential to ensure that the opinions expressed are correctly attributed to the right person.
Thus, low precision associated with hallucinations undermines the credibility of the system. On the other hand, low precision without hallucinations is not so harmful: as long as the system has sufficiently summarized the transcript, it provides the user with a broad range of opinions to write a news article.
@@ 4.1.2 Hallucination
To check for hallucination in the summarization process of a transcript, we consider the following steps:
@@
    For each person, split their speech into small chunks. In this article, the text was split into chunks of 5 sentences (with an overlap of 1 sentence) using the spaCy [^8] https://spacy.io/ framework with the pt_core_news_sm model;
@@
    Create a vector database for each person using the chunks. In this article, we used the OpenAI’s text-embedding-3-small model to encode the chunks;
@@
    For each opinion returned in the structured summary of the experiment, search the vector database of the corresponding individual to find the most relevant chunks related to that person’s opinion. The opinions were also encoded using the OpenAI’s text-embedding-3-small model;
@@
    Send the opinion returned in the summarization and the four most relevant chunks found in the previous step to a LLM to check if the generated information can be inferred from the chunks. Three prompts, described in section 4.2, were tested in this step.
Using this strategy, it is possible to summarize a transcript to extract the opinions of the individuals participating in the hearing and to check if the summarization system is capturing most of the relevant opinions (recall). It also allows us to check if the system is generating many opinions that are not as important (precision). Finally, it is possible to identify and remove most of the hallucinations.
@@ 4.2 Results and discussion
In this experiment, we prompted ChatGPT using the custom GPT to generate a metadata structure similar to the one provided by the dataset, which includes individuals and their opinions. Considering all the transcripts in the dataset, 4 238 opinions were extracted by the custom GPT.
Table 2 shows the statistics of recall and precision using the proposed method after running the experiment in the 206 samples of the dataset. The average recall was 44.9%, meaning the custom GPT prompt recovered nearly half of the relevant opinions for each sample of the dataset. On the other hand, the average precision was 24.8%, indicating that for each relevant opinion in the sample, the custom GPT also generated three other opinions.
@@
To check the level of hallucination in the generated opinions, we initially applied the first four steps described in the method presented in section 4.1 to retrieve the four most similar chunks for each opinion and manually annotated the 4 238 opinions generated in the experiment [^9] The data generated in this operation was also provided in PublicHearingBR dataset, allowing it to be used in NLI tasks in Portuguese.. This was done by reading the chunks and checking whether the opinion could be inferred from them or not. If it was not possible, it was annotated as a hallucination.
@@
The manual annotation showed that 3 734 opinions can be inferred from the four closest chunks, while 504 (11.89%) cannot be inferred and were considered hallucinations. This value is an upper limit of the actual number of hallucinations generated by the custom GPT, as it is possible that the opinion is supported by text that was not in the four closest chunks.
@@
After the manual annotation, we applied the final step of the method to detect hallucinations. We tested three prompts (Figure 7, 8, and 9) using the GPT-4o mini and the GPT-4o models to automatically identify whether the opinion can be inferred from the chunks. All prompts request the same information but ask for it in different ways.
@@
The first prompt asks if the opinion can be completely inferred from the chunks and requests that the answer be given in JSON format with the reasoning for the decision.
@@
The second prompt is an evolution of the first one and explicitly requests sentences that support the decision before giving the reasoning. Besides, it states that the model is an expert in text analysis.
@@
The third prompt evolves from the second one by reinforcing the request and stating that, if in doubt, it should answer that the opinion cannot be inferred from the chunks. The prompt also defines that the model is an expert in discourse analysis.
Tables 3, 4, and 5 show the confusion matrix for the three prompts tested for automatic hallucination detection using the GPT-4o mini model. Tables 6, 7, and 8 shows the same information using the GPT-4o model. Regardless of the tested prompt, the proposed method for detecting hallucinations identify most of the hallucinations. However, there is a clear trade-off between the percentage of correctly detected hallucinations and the number of valid opinions that were identified as hallucinations (these false positives will be referred to as false hallucinations in this article).
Table 3: Confusion matrix for automatic hallucination detection using Prompt 1 and GPT-4o mini.
Prompt 1
Can be inferred Hallucination Total
Dataset Can be inferred 3 673 61 3 734
Hallucination 181 323 504
Total 3 854 384 4 238
Table 4: Confusion matrix for automatic hallucination detection using Prompt 2 and GPT-4o mini.
Prompt 2
Can be inferred Hallucination Total
Dataset Can be inferred 3 274 460 3 734
Hallucination 89 415 504
Total 3 363 875 4 238
Table 5: Confusion matrix for automatic hallucination detection using Prompt 3 and GPT-4o mini.
Prompt 3
Can be inferred Hallucination Total
Dataset Can be inferred 2 405 1 329 3 734
Hallucination 39 465 504
Total 2 444 1 974 4 238
Table 6: Confusion matrix for automatic hallucination detection using Prompt 1 and GPT-4o.
Prompt 1
Can be inferred Hallucination Total
Dataset Can be inferred 3 561 173 3 734
Hallucination 165 339 504
Total 3 726 512 4 238
Table 7: Confusion matrix for automatic hallucination detection using Prompt 2 and GPT-4o.
Prompt 2
Can be inferred Hallucination Total
Dataset Can be inferred 3 276 458 3 734
Hallucination 90 414 504
Total 3 366 872 4 238
Table 8: Confusion matrix for automatic hallucination detection using Prompt 3 and GPT-4o.
Prompt 3
Can be inferred Hallucination Total
Dataset Can be inferred 2 922 812 3 734
Hallucination 46 458 504
Total 2 968 1 270 4 238
For the GPT-4o mini model, Prompts 1, 2, and 3 correctly detected, respectively, 64.09%, 82.34%, and 92.26% of the hallucinations in the dataset. There is a trade-off between the detection of hallucinations and the number of false hallucinations (valid opinions identified as hallucinations). The false hallucinations rates were, respectively, 1.63%, 12.32%, and 35.59%, for prompts 1, 2, and 3.
For the GPT-4o model, Prompts 1, 2, and 3 detected, respectively, 67.26%, 82.12%, and 90.87% of the hallucinations in the dataset. The false hallucinations rates were, respectively, 4.63%, 12.27%, and 21.75%.
Figure 2 shows this data. Regarding the percentage of correctly detected hallucinations, considering identical prompts, there was a small difference between the models: 3.17 percentage points for Prompt 1 (best result with GPT-4o), 0.22 percentage points for Prompt 2 (best result with GPT-4o mini), and 1.39 percentage points for Prompt 3 (best result with GPT-4o mini).
However, we observed a better trade-off between the percentage of correctly detected hallucinations and the number of false hallucinations with the GPT-4o model. This can be seen when we compare Prompt 2 to Prompt 3: in this case, the GPT-4o mini model showed an increase in the percentage of correctly detected hallucinations of 9.92 percentage points, but this came at the cost of a 23.27 percentage point increase in the percentage of false hallucinations. In contrast, for the GPT-4o model, the increase in the percentage of correctly detected hallucinations of 8.75 percentage points came with a much smaller cost of a 9.48 percentage point increase in the percentage of false hallucinations.
@@
Figure 2: (Left) percentage of correctly detected hallucinations and (right) false hallucinations for the three prompts tested with the GPT-4o and GPT-4o mini models.
Nevertheless, we understand that the proposed method is a valid approach to detect hallucinations. For Prompt 3 and the GPT-4o model (Table 8), even though 812 opinions were classified as hallucinations, there were still 2 922 useful opinions left (about 14 opinions per transcript). Only 46 out of 504 (9.13%) hallucinations went undetected. This number should be seen as the maximum number of hallucinations present, as the classification of an opinion as a hallucination was made considering four chunks (rather than the entire transcript).
The comparison with the different prompts and the GPT-4o and GPT-4o mini models showed that the result depends on both the quality of the prompt and the LLM used. Thus, the results suggest that, with the advancements of LLMs and the application of more sophisticated prompt engineering techniques, the trade-off between the percentage of correctly detected hallucinations and the number of false hallucinations should improve. Moreover, we understand that the results also rely on the retrieval stage of the chunks that will be sent to the model. Hence, different configurations (for example, changing the number of the chunks or using a better embedding model) may yield better results. However, due to the statistical nature of the methods used, we currently do not foresee a way to solve this without any trade-off.
@@
It is worth noting that, in this experiment, the entire transcript was summarized all at once, simulating the perspective of an end user of ChatGPT. In this case, two types of hallucinations were generated in the experiment. The first type is when an opinion expressed by one person is attributed to another, e.g., attributing to Bob an opinion that was voiced by Alice. The second type of hallucination is when an incorrect text is generated by the LLM. We believe that in a real summarization system, it would be ideal to first separate each individual’s speech and then summarize each speech separately. In this case, we can remove hallucinations of the first type (attributing a valid opinion to the wrong person), possibly reducing the total number of hallucinations.
@@ 5 Conclusion
We introduced PublicHearingBR, a new Portuguese dataset with 206 samples for the summarization of long documents (transcripts of public hearings held by the Brazilian Chamber of Deputies). We also presented, as a baseline, a custom GPT prompt to summarize the transcripts, extracting the main individuals of the hearing and their opinions. We discussed evaluation metrics for this task, with emphasis on detecting hallucinations in the generated summaries. The discussion on hallucination resulted in 4 238 manually annotated samples that can be used to test Natural Language Inference (NLI) systems in Portuguese.
There is a lack of summarization datasets in Portuguese, and the existing ones for summarizing large documents do not include a summary of the main discussions carried out in the document. PublicHearingBR, in addition to addressing this issue, contains summaries that have been manually verified.
We experiment the dataset with a custom GPT prompt to summarize public hearings. The goal is not to solve this problem, but rather to use a tool available to end users to establish a baseline that is useful for future studies.
In this experiment to extract people’s opinions, on average almost 45% of the opinions of the samples in the dataset was extracted (recall). Additionally, the average precision was approximately 25% per sample, indicating that it returned three times more opinions that were not in the reference dataset. We manually verified the 4 238 generated opinions and checked whether they could be inferred with the four closest chunks of text. This analysis indicates that almost 12% of the generated opinions could be hallucinations.
We proposed a method for verifying hallucinations that relies on using an LLM as a judge. We found that there is a clear trade-off between the number of correctly detected hallucinations and the number of false hallucinations identified. The more hallucinations are identified, the greater the number of false hallucinations.
The experiment conducted, although yielding interesting results, indicates that a real summarization system needs to be more complex than what current off-the-shelf systems allow.
Thus, as a possible direction for future research, we identified the design of a public hearing summarization system, where summarization is performed in parts, initially splitting the speeches of each individual before summarizing, in order to reduce the probability of hallucinations. Other possibilities for future research include using the dataset for the automatic generation of a news article based on the metadata or the public hearing’s transcript itself.

### v2

@@ 4.1 Method: Description and evaluation of the experiment
This experiment addresses the primary use case of the dataset, as described in Section 3.1: summarizing long documents to extract the most relevant opinions of the individuals.
As shown in Section 3.3, the transcripts in the dataset are long documents and thus can exceed the context window of some LLMs. Instead of building a custom pipeline to handle this, we tested a general-purpose, off-the-shelf generative AI tool: ChatGPT, given its widespread use. This approach simulates how a typical user might handle the task with general-purpose software and also establishes a baseline for the dataset, which can be improved as more specialized systems become available.
We used ChatGPT with a instruction prompt (Figure 5) that guided the model on how to read and interpret the transcripts. The prompt specified the input data format and instructed the language model to extract a structured summary of the most relevant opinions from the transcript.
Once ChatGPT has extracted the structured summaries for all samples, they need to be compared with the ground truth provided in the PublicHearingBR dataset to assess the quality of the summarization.
Since this summarization approach is hybrid, using the ROUGE metric [5] is inadequate as it relies on n-grams rather than semantic content. We propose that the evaluation covers three aspects: 1. recall: to measure the percentage of relevant opinions the experiment was able to retrieve; 2. precision: to determine the percentage of retrieved opinions that are relevant; 3. hallucination: to ensure that the extracted opinions are present in the transcript, even if they are not in the dataset.
@@ 4.1.1 Recall and precision
We use the following steps to calculate recall and precision:
@@
    First, individuals mentioned in the transcript are mapped to those in the metadata available in PublicHearingBR. Since names might differ (due to abbreviations or titles, for example), a language model is initially used to perform this mapping. This was done using GPT-4o with the prompt shown in Figure 6.
@@
    Next, for each individual identified in the previous step, the opinions extracted from the transcript are mapped to those in the PublicHearingBR dataset. Because these are natural language texts, it is important to determine whether opinions are semantically equivalent. This mapping was also performed using GPT-4o with the prompt shown in Figure 7.
@@
    Once the mapping is complete, the number of matching opinions between the extracted data and the dataset can be counted, allowing the calculation of recall and precision.
Recall is a relevant metric because it measures the proportion of reference opinions – identified by human annotators in the dataset – that are successfully retrieved by the summarization system. It is worth noting that the recall calculated through the mapping performed by the LLM is the essence of the Check-Eval method. The main difference is that Check-Eval automatically generates the reference checklist from a text using an LLM [6], while this method obtains the checklist directly from the dataset.
Precision is another important metric to consider, although low precision is not necessarily a negative outcome. For example, a journalist might omit some relevant opinions due to space constraints. However, it is important to assess if the system is hallucinating. This is especially important because the goal is to write a news article (or to generate inputs for one), where false information is unacceptable. Besides, since participants in public hearings share their ideas and opinions, the opinions must be attributed to the correct individuals.
Thus, low precision associated with hallucinations undermines the credibility of the system. On the other hand, low precision without hallucinations is less problematic: as long as the system sufficiently summarizes the transcript, it still offers the user a broad range of opinions to write a news article.
@@ 4.1.2 Hallucination
To detect hallucinations in the summarization of a transcript, we follow the steps below:
@@
    Split each person’s speech into small chunks. In this article, the text was split into chunks of five sentences with a one-sentence overlap using the spaCy [^9] https://spacy.io/ framework and the pt_core_news_sm model;
@@
    Create a vector database for each person using these chunks. We used OpenAI’s text-embedding-3-small model to encode the chunks;
@@
    For each opinion in the structured summary produced by the experiment, search the corresponding individual’s vector database to find the most relevant chunks. Opinions were also encoded using OpenAI’s text-embedding-3-small model;
@@
    Send the opinion extracted by the summarization and the four most relevant chunks to an LLM to check if the generated information can be inferred from the chunks. Three prompts, detailed in Section 4.2, were tested in this step.
Using this strategy, it is possible to summarize a transcript by extracting the opinions of the individuals participating in the hearing and to check if the summarization system captures most of the relevant opinions (recall). It also allows us to check if the system generates many opinions that are not relevant (precision). Finally, it is possible to identify and remove most of the hallucinations.
@@ 4.2 Results and discussion
4.2.1 Precision and recall
In this experiment, we prompted ChatGPT to generate a structured summary (metadata) that includes individuals and their opinions. Considering all the transcripts in the dataset, a total of 4 238 opinions were extracted.
Table 2 shows the recall and precision statistics from our experiment on the 206 dataset samples. The average recall was 44.9%, meaning the ChatGPT retrieved nearly half of the relevant opinions for each sample. On the other hand, the average precision was 24.8%, which indicates that, for every relevant opinion generated, it also produced roughly three other opinions.
@@ 4.2.2 Hallucination
To assess the level of hallucination in the generated opinions, we first manually annotated the generated content to identify hallucinations, in order to validate the automatic annotation method. To do so, we applied the first four steps described in Section 4.1.2 to retrieve the four most similar text chunks for each of the 4 238 opinions generated in the experiment. A consultant from the Chamber of Deputies then reviewed each opinion and its four corresponding chunks, indicating whether the opinion could be inferred from them. If so, the opinion-chunks pair was annotated as a valid opinion; otherwise, as a possible hallucination.
@@
The manual annotation showed that 3 734 opinions are valid (i.e., can be inferred from the four closest chunks), while 504 (11.89%) are possible hallucinations. This percentage represents an upper bound and may be an overestimate of the actual number of hallucinations generated by the experiment, as some opinions might be supported by text outside the four closest chunks. From now on, “possible hallucinations” will be referred to simply as “hallucinations”.
We understand that this manual annotation of 4 238 samples indicating whether an opinion can be inferred from text excerpts can also be used to evaluate NLI systems in Portuguese. Therefore, we also provide this data as a JSONL file [^10] PublicHearingBR_NLI.jsonl, available in the repository of this article..
Automatic annotation
@@
After the manual annotation, we applied the final step of the method to detect hallucinations: we used an LLM to check if an opinion could be inferred from the four closest chunks of text. We tested three prompts with four LLMs: GPT-4o mini, GPT-4o, DeepSeek-V3, and Sabiá-3.1 (a Brazilian Portuguese model). All prompts request the same information but phrase the request differently:
@@
  • Prompt 1 (Figure 8) asks if the opinion can be fully inferred from the chunks and requests the answer in JSON format, including the reasoning behind the decision.
@@
  • Prompt 2 (Figure 9) builds on Prompt 1 and explicitly requests the sentences that support the decision before providing the reasoning. It also states that the model is an expert in text analysis.
@@
  • Prompt 3 (Figure 10) extends Prompt 2 by reinforcing the request and specifying that, in case of doubt, the model should conclude that the opinion cannot be inferred from the chunks. It also defines the model as an expert in discourse analysis.
Results and discussion of the automatic hallucination evaluation
For each of the three prompts and four LLMs tested for automatic hallucination evaluation, we generated a confusion matrix comparing the output with the 4 238 manually annotated opinions (3 734 valid opinions and 504 hallucinations). Tables 5 to 8 in Appendix C show the matrices.
Tables 3 and 4 provide an overview of each model and prompt’s performance in terms of correctly detected hallucinations (true positives) and valid opinions labeled as hallucinations (false positives). Figure 2 shows the same information, including a 95% confidence interval. Finally, Figure 3 illustrates the trade-off between the percentage of true positives and false positives across models and prompts.
Table 3: Hallucinations correctly identified by models and prompts (true positives). Percentages are based on the 504 manually annotated hallucinations.
Model Prompt 1 Prompt 2 Prompt 3
GPT-4o mini 323 (64.09%) 415 (82.34%) 465 (92.26%)
GPT-4o 339 (67.26%) 414 (82.12%) 458 (90.87%)
DeepSeek-V3 263 (52.18%) 337 (66.86%) 409 (81.15%)
Sabiá-3.1 249 (49.4%) 283 (56.15%) 296 (58.73%)
Table 4: Valid opinions incorrectly identified as hallucinations by models and prompts (false positives). Percentages are based on the 3 732 manually annotated valid opinions.
Model Prompt 1 Prompt 2 Prompt 3
GPT-4o mini 61 (1.63%) 460 (12.32%) 1 329 (35.59%)
GPT-4o 173 (4.63%) 458 (12.27%) 812 (21.75%)
DeepSeek-V3 68 (1.82%) 149 (3.99%) 434 (11.62%)
Sabiá-3.1 57 (1.54%) 87 (2.33%) 102 (2.73%)
@@
Figure 2: (Left) percentage of correctly detected hallucinations and (right) false hallucinations for the three prompts tested with GPT-4o mini, GPT-4o, DeepSeek-V3, and Sabiá-3.1 models.
Figure 3: Trade-off between hallucinations detected (true positives) and valid opinions tagged as hallucinations (false positives) across models and prompts. The dashed lines show the linear regression for each model (colors) and across all points (black).
The first point to observe is the effect of prompt engineering. As the prompts became more precise – explicitly requiring evidence, emphasizing direct inference, or including strict directives – the true positive rate increased. Consequently, for a given LLM, Prompt 1 detected the fewest hallucinations, followed by Prompt 2 and then Prompt 3.
However, as illustrated in Figure 3, there is a classic trade-off between the true positive rate (correctly hallucinations detected) and the false positive rate (valid opinions labeled as hallucinations). Models and prompts that detect more hallucinations tend to make more errors when classifying valid opinions.
Regardless of the model tested, increasing the total number of detected hallucinations exponentially increases the total number of valid opinions labeled as hallucinations. For the GPT-4o, DeepSeek-V3, and Sabiá-3.1 models, this increase follows a similar exponential pattern ($\approx e^{0.06x}$, where $x$ is the percentage of true positives). However, for the GPT-4o mini model, the growth is steeper ($\approx e^{0.11x}$). In this trade-off, GPT-4o mini performed the worst, while the other three models showed comparable behavior.
Considering the models with similar trade-off behavior and the three prompts tested, the best performance was observed with GPT-4o, followed by DeepSeek-V3 and Sabiá-3.1. However, since the trade-off between true positives and false positives is similar across these models, it is possible that other prompt variations would yield different overall results without changing this underlying relationship. Therefore, this information alone is not sufficient to conclude that GPT-4o would outperform the others with a different set of prompts.
Regarding the susceptibility of the models to prompt variation, the Sabiá-3.1 model was the least affected: Prompt 3 detected fewer than 10 percentage points more hallucinations than Prompt 1. In contrast, GPT-4o mini was the most susceptible, with an increase of approximately 28 percentage points in hallucination detection from Prompt 1 to Prompt 3 (at the cost of an increase of about 34 percentage points in the percentage of false positives).
The proposed method is a valid way to detect hallucinations, but its effectiveness depends on the LLM and prompt used. With Prompt 3 and the GPT-4o model (Table 6), 812 valid opinions were misclassified as hallucinations, while 2 922 valid opinions remained (about 14 per transcript). Among the 504 hallucinations in the dataset, only 46 went undetected (9.13%). In other words, the method labeled 2 968 opinions as valid (2 922 correct + 46 missed hallucinations), meaning that about 1.5% of the estimated valid opinions may still be hallucinations. This value should be seen as an upper bound, since the method to detect hallucinations used only four chunks of text (i.e., some of these 46 missed hallucinations might be valid opinions supported by text outside these chunks).
The results were obtained for 12 configurations (combinations of 3 prompts and 4 models). The trade-off relationships between true and false positives found were similar in 3 of the 4 models. This suggests that the relationship is not fixed; in other words, advances in LLMs may reduce this number, improving the use of LLMs for this task.
In addition, the automatic hallucination detection in this experiment did not account for the effect of the chunks retrieved and sent to the LLMs. Performance also depends on the chunk retrieval strategy. Adjustments such as increasing the number of retrieved chunks or using more accurate embedding models could lead to better results. Nevertheless, given the probabilistic nature of the methods, it is unlikely that this trade-off can be eliminated entirely.
@@
It is worth noting that, in this experiment, each transcript was summarized as a whole, simulating the perspective of a typical ChatGPT user. In this scenario, we identified two types of hallucinations. The first occurs when an opinion expressed by one person is attributed to the wrong speaker: for example, attributing to Bob an opinion actually voiced by Alice. The second involves factually incorrect content. In a dedicated system for this task, a better approach would be to isolate each participant’s speech and summarize it individually. This could eliminate the first type of hallucination (misattributed opinions) and likely reduce the overall number of hallucinations.
@@ 5 Conclusion
We introduced PublicHearingBR, a new Portuguese dataset with 206 samples for summarizing long documents (transcripts of public hearings held by the Brazilian Chamber of Deputies). We also presented a baseline using ChatGPT to extract the main individuals involved in the hearings and their opinions from the transcripts. Furthermore, we discussed evaluation metrics for this task, with a focus on detecting hallucinations in the generated summaries. This discussion resulted in 4 238 manually annotated samples, which can be used to evaluate natural language inference systems in Portuguese.
There is a lack of summarization datasets in Portuguese, and the existing ones for long-document summarization do not include summaries of the main discussions presented in source texts. PublicHearingBR addresses this gap and, in addition, provides summaries that have been manually verified.
We experimented with the dataset using ChatGPT and a custom prompt to summarize public hearings. The goal is not to solve the task, but to establish a baseline using a tool readily available to end users, supporting future research.
In the experiment, which focused on extracting individuals’ opinions, ChatGPT achieved an average recall of nearly 45%, meaning it retrieved almost half of the opinions present in the reference dataset. The average precision was approximately 25% per sample: for each relevant opinion, the system generated three others not present in the reference. We manually verified the 4 238 generated opinions and checked whether they could be inferred from the four closest chunks of text. This analysis indicated that nearly 12% of the generated opinions may be hallucinations.
We proposed a method for hallucination detection that uses an LLM as a judge. Our findings reveal a clear trade-off: the more hallucinations are detected, the more valid opinions are labeled as hallucinations. We analyzed this relationship across three prompts and found that, for the GPT-4o, DeepSeek-V3, and Sabiá-3.1 models, the increase in valid opinions labeled as hallucinations follows a similar exponential pattern ($\approx e^{0.06x}$, where $x$ is the percentage of hallucinations detected). In contrast, for the GPT-4o mini model, the growth is steeper: $\approx e^{0.11x}$.
The experiment yielded relevant insights and suggests that a real-world summarization system needs to be more complex than current off-the-shelf tools allow. As a direction for future research, we propose designing a dedicated summarization system for public hearings where the process is divided into stages, starting with the separation of each participant’s speech to reduce the risk of hallucinations. Other possibilities include using the dataset to automatically generate news articles based on either the metadata or the full transcript of the hearing.